OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (anonymous):

A. 3x - 3y = -8 B. x - 3y = -9 C. x - 3y = -8 D. 3x + y = 9